# Play2Prompt: Zero-shot Tool Instruction Optimization for LLM Agents via Tool Play

### v1

@@ Abstract
Large language models (LLMs) are increasingly integrated with specialized external tools, yet many tasks demand zero-shot tool usage with minimal or noisy documentation. Existing solutions rely on manual rewriting or labeled data for validation, making them inapplicable in true zero-shot settings. To address these challenges, we propose Play2Prompt, an automated framework that systematically “plays” with each tool to explore its input-output behaviors. Through this iterative trial-and-error process, Play2Prompt refines tool documentation and generates usage examples without any labeled data. These examples not only guide LLM inference but also serve as validation to further enhance tool utilization. Extensive experiments on real-world tasks demonstrate that Play2Prompt significantly improves zero-shot tool performance across both open and closed models, offering a scalable and effective solution for domain-specific tool integration [^1] Source code available at https://github.com/wfangtw/play2prompt.
@@ 1 Introduction
We evaluate Play2Prompt with benchmark on real-world scenarios. On the Berkeley Function-Calling Leaderboard (Yan et al., 2024) and the StableToolBench benchmark (Guo et al., 2024), our approach consistently surpasses baseline methods for both open LLaMA models (Dubey et al., 2024a) and closed OpenAI GPT models (Achiam et al., 2023). Extensive experiments and analyses further underscore the robustness of our approach.
@@ 3.1 Experimental Setup
ReAct+EasyTool 49.9 69.0 65.2 53.8 62.1 59.0 61.1
@@
Dataset: StableToolBench.
LLaMA-8B Prompting 91.0 67.1 92.0 84.0 70.0 81.3 80.8
+Play2Prompt 94.0 81.4 92.0 82.0 82.5 86.1 86.4
LLaMA-70B Prompting 96.0 90.0 92.0 80.0 72.5 84.4 86.1
+Play2Prompt 95.0 88.6 92.0 84.0 87.5 88.8 89.4
GPT-3.5 Function-calling 92.0 87.1 84.0 76.0 50.0 74.9 77.8
+Play2Prompt 98.0 88.6 86.0 84.0 80.0 85.8 87.3
GPT-4o Function-calling 93.0 95.7 92.0 88.0 75.0 87.3 88.7
+Play2Prompt 98.0 94.3 94.0 88.0 82.5 90.3 91.6
@@
Inference and Evaluation.
We adhere to the official inference configurations of each benchmark, where a set of tools is provided for each test query. For task LLM ${\mathcal{M}}_{T}$, we tested llama-3.1-8b-instruct and llama-3.3-70b-instruct for LLaMA (Dubey et al., 2024b), while gpt-3.5-turbo-0125 and gpt-4o-2024-11-20 were used for GPT (Achiam et al., 2023). For BFCL, single-turn prompting with official prompts is used as the baseline inference method for LLaMA models, while direct function-calling mode is used for the GPT models. We do not provide tool-use examples or additional documentation in these baseline runs, complying with a zero-shot setting. Play2Prompt instead applies optimized in-context examples and documentation as prompts and runs the baseline inference methods. It supports multi-tool queries by independently optimizing examples and documentation for each tool and then performing inference. The official evaluation metric of accuracy is used, using exact matches or structural matches (depending on categories) to determine correctness. We report the average across five categories, and, following the official setup, the weighted average (“Simple” categories weighted by $0.5$).
StableToolBench employs a more complex chain-of-thought inference method, ReAct (Yao et al., 2023). We also compare against EasyTool Yuan et al. (2024), which uses direct prompting to optimize tool documentation and generate usage scenarios, but requires documentation and labeled in-context examples from additional tools and is thus not entirely zero-shot [^3] We use the optimized tool documentation and usage scenarios provided at https://github.com/microsoft/JARVIS/tree/main/easytool. The inference setting of Yuan et al. (2024) differs from ours as we follow the official inference prompts provided with StableToolBench.. An evaluation LLM is used to judge whether a response adequately answers a user query, due the dataset’s free-form output design. We follow the official pipeline, using official ReAct prompts and report solvable pass rate, which measures the percentage of queries deemed solvable by the evaluation LLM. We use llama-3.3-70b-instruct as the evaluation LLM, which produces stable assessments. Further details are provided in appendix C.
Optimization Details.
Play2Prompt first uses beam search to optimize tool-use examples. We set $\lambda=1$, $N_{E}=3$, limit depth to $3$, and explore $L=3$ beams per node. We use beam width $W=10$ for BFCL and $W=3$ for StableToolBench to generate and select the top $W$ examples for each tool, which then are then passed to the documentation optimization procedure. Beam search is again applied to select the best tool documentation. We employ llama-3.1-8b-instruct as both ${\mathcal{G}}_{E}$ and ${\mathcal{G}}_{D}$.
3.2 Results and Analyses
Results on BFCL and StableToolbench.
In table 1, we summarize the results on BFCL with LLaMA and GPT task models, both with and without tool-use examples and tool documentation produced by Play2Prompt. With Play2Prompt, absolute gains of 5-7% are observed for LLaMA models and GPT-3.5, while GPT-4o achieves increases of 2.5% and 3.3% in average and weighted average accuracy, respectively, despite its already high baseline. Notably, Play2Prompt addresses challenging categories such as REST for LLaMA-8B, and Multiple-Parallel for all models, yielding improvements of 12-17%. These gains suggest that optimized in-context examples and documentation can correct specific shortcomings in tool usage.
In table 2, we show the solvable pass rates on StableToolBench for LLaMA and GPT models with and without Play2Prompt. Our approach surpasses all baselines with average gains of 5-7% across LLaMA models and GPT-3.5, and outperforms the specifically designed EasyTool, which is not fully zero-shot as it requires in-context demonstrations during optimization. Play2Prompt is also consistent across subsets, avoiding large performance drops in multiple subsets when using EasyTool, highlighting the benefits of real-time “tool-play” and evaluation of candidate examples and documentation during optimization. Moreover, as GPT-4o exhibits more sensitivity to documentation and in-context demonstration changes on this dataset, EasyTool degrades GPT-4o on all subsets, whereas Play2Prompt achieves a 2% improvement on average. Overall, Play2Prompt essentially boosts performance of models up to the baseline performance of the larger models. It is noteworthy that most test samples for all 6 subsets in StableToolBench, as well as the Multiple and Multiple-Parallel subsets in BFCL, are multi-tool queries (see Section 3.1), underscoring the effectiveness of Play2Prompt, which operates on single tools independently during optimization.
@@ 3.2 Results and Analyses
To assess the robustness of Play2Prompt, we artificially introduce noise to tool documentation in BFCL Executable by dropping each parameter description with a probability $p=0.5$, leaving general tool descriptions and signatures intact. As shown in table 3, this degradation reduces baseline performance by about 5% for the LLaMA models and GPT-4o, and by 10% for GPT-3.5, especially for the most challenging Multiple-Parallel category. With Play2Prompt, performance consistently exceeds baselines across all models, improving by 3-4% for the larger LLaMA-70B and GPT-4o models, 5% for LLaMA-8B, and most significantly, recovering the full 10% loss for GPT-3.5. These results illustrate Play2Prompt’s robustness to incomplete documentation and highlight its potential in lower-resource settings.
@@
Search strategy I3-Inst
ReAct 64.3
+P2P-MC (depth$=1$) 65.2
+P2P-MC (depth$=5$) 66.7
+P2P-MC (depth$=5$) & $N_{E}$$=3$ 68.9
+P2P-BeamSearch (depth$=5$) & $N_{E}$$=3$ 70.7
@@
We investigate alternative strategies for exploring the example-documentation space by comparing beam search in Play2Prompt to Monte Carlo search (MC) at various depths, sharing the same sample proposal scheme but with different number of proposals and sub-sampling approach.
We also study the role of $N_{E}$ self-reflection steps during example generation. Results are shown in table 6, focusing on I3-Inst, which combines multiple tools from different categories and is thus more challenging. The results show that MC underperforms beam search due to its limited exploration and is more subject to randomness, while deepening the search and adding self-reflection actions significantly improves performance, highlighting the importance of systematic search in optimizing examples and documentation.
@@ 4 Related Work
LLMs for Tool Use
Recent years have seen notable advances in employing large language models (LLMs) as agents to master tool use for solving complex tasks (Mialon et al., 2023; Qin et al., 2024a), thereby enhancing LLMs’ capabilities in multi-modal understanding (Gupta and Kembhavi, 2023; Surís et al., 2023; Wu et al., 2023), programming tools (Gao et al., 2023; Paranjape et al., 2023; Team et al., 2023; Zhang et al., 2023b; Cai et al., 2024), and other domain-specific functionalities. The conventional approach involves training base models with tool-use data (Thoppilan et al., 2022; Dubey et al., 2024a) or fine-tuning LLMs (Patil et al., 2023; Schick et al., 2023; Yang et al., 2023; Parisi et al., 2022) on specific tasks with a small number of tools, but requires continual learning as new tools are added. Hao et al. (2023) addressed scalability by training tool embeddings plug-and-play usage, though still requiring labeled data. Alternatively, LLMs can be augmented with meta-prompts or tool-use instructions at inference time (Lu et al., 2023; Shen et al., 2023; Song et al., 2023; Qin et al., 2024b; Zhuang et al., 2024). As the range of applications and tools expands, enhancing LLMs’ capacity to handle new tools remain pivotal, which we address through Play2Prompt.
Tool-Use Instructions and Optimization
Tool documentation and example demonstrations are key to prompting LLMs for effective tool use. Hsieh et al. (2023) highlighted the risk of hallucinations when documentation is lacking, and Xu et al. (2023) showed performance declines without in-context examples. To automate generating tool-use instances, Shen et al. (2024) leveraged a graph of tool relations for back-instructing queries, relying on the availability of external tool graphs. Yuan et al. (2024) proposed direct prompting to rewrite tool documentation, which relies on labeled documentation examples and lacks systematic optimization and the ability to measure optimization quality. Automatic prompt tuning (Pryzant et al., 2023; Wang et al., 2024) adapts prompts to domain-specific tasks but require held-out test sets, rendering it unsuitable for zero-shot tool instruction rewriting (Wu et al., 2024). These constraints underscore the need for approaches that optimize tool instructions and demonstrations without labeled data or manual effort, which Play2Prompt achieves by interacting directly with the tool itself.
@@ 5 Conclusion
In this work, we generate proxy testing sets based only on a single given tool and do not cover multi-tool use. Scaling from single-tool scenarios to multiple tools can likely enhance LLM’s tool use effectiveness. Additionally, for example demonstrations, we use rejection sampling to generate tool invocations first, which do not work for functions whose parameter space is too large, for instance parameters that take long ID string inputs or authentication tokens that require calls to other tools beforehand. Exploring multi-tool dependencies could potentially resolve this issue and improve tool play. In our work we focus on tool documentation and demonstrations only, relegating other information as meta-information, which could be potential next steps to explore.
@@ Appendix A Dataset License
Appendix B Detailed Optimization Procedures
Appendix C Additional Implementation Details
For GPT models, we keep the exact same inference setting as used in the original benchmark, except for requiring it to return a single action at each step from the provided toolset by supplying a flag to the OpenAI API. For LLaMA models, on BFCL we follow the exact official inference prompts, and on StableToolBench we adapt the ReAct prompts into LLaMA-3’s prompt format but keep everything else fixed as much as possible.

### v2

@@ Abstract
Large language models (LLMs) are increasingly integrated with specialized external tools, yet many tasks demand zero-shot tool usage with minimal or noisy documentation. Existing solutions rely on manual rewriting or labeled data for validation, making them inapplicable in true zero-shot settings. To address these challenges, we propose Play2Prompt, an automated framework that systematically “plays” with each tool to explore its input-output behaviors. Through this iterative trial-and-error process, Play2Prompt refines tool documentation and generates usage examples without any labeled data. These examples not only guide LLM inference but also serve as validation to further enhance tool utilization. Extensive experiments on real-world tasks demonstrate that Play2Prompt significantly improves zero-shot tool performance across both open and closed models, offering a scalable and effective solution for domain-specific tool integration [^1] Source code is available at https://github.com/wfangtw/play2prompt..
@@ 1 Introduction
We evaluate Play2Prompt with benchmark on real-world scenarios. On the Berkeley Function-Calling Leaderboard (Yan et al., 2024) and the StableToolBench benchmark (Guo et al., 2024), our approach consistently surpasses baseline methods for both open (Dubey et al., 2024; Liu et al., 2025; Lin et al., 2025) and closed models (Achiam et al., 2023). Extensive experiments and analyses further underscore the robustness of our approach.
@@ 3.1 Experimental Setup
ToolACE-8B Function-calling 96.0 92.9 92.0 86.0 70.0 85.6 87.4
+Play2Prompt 95.0 90.0 94.0 90.0 82.5 89.8 90.3
Hammer-7B Function-calling 96.0 72.9 88.0 86.0 70.0 82.1 82.6
+Play2Prompt 95.0 82.8 92.0 86.0 80.0 86.7 87.2
@@
ReAct+EasyTool 49.9 69.0 65.2 61.3 65.3 52.5 60.5
ReAct+DRAFT - - - - 62.1 63.9 -
@@
Inference and Evaluation.
We adhere to official inference settings of each benchmark, where a set of tools is provided for each test query. For task LLM ${\mathcal{M}}_{T}$, we tested the 8B and 70B LLaMA models (Dubey et al., 2024), and GPT-3.5 and GPT-4o were used for GPT (Achiam et al., 2023). We also tested on BFCL two state-of-the-art LLMs trained specifically for tool-calling, namely ToolACE (Liu et al., 2025) and Hammer (Lin et al., 2025), with both models achieving high rankings on the BFCL leaderboard as of this publication. For BFCL, single-turn prompting with official prompts is used as the baseline inference method for LLaMA models, while direct function-calling mode is used for the GPT models, ToolACE, and Hammer. We do not provide tool-use examples or additional documentation in these baseline runs, complying with a zero-shot setting. For Play2Prompt, optimized in-context examples and documentation are supplied as prompts and runs the baseline inference methods. It supports multi-tool queries by independently optimizing examples and documentation for each tool and then performing inference. We use the official evaluation metric of accuracy, with exact or structural matches depending on categories to determine correctness. We report the average across five categories, and, following the official setup, the weighted average (“Simple” categories weighted by $0.5$).
StableToolBench employs a more complex chain-of-thought inference method, ReAct (Yao et al., 2023). We compare against EasyTool Yuan et al. (2024), which uses direct prompting to optimize tool documentation and generate usage scenarios, but requires documentation and labeled in-context examples from additional tools and is thus not entirely zero-shot [^3] We use the optimized tool documentation and usage scenarios provided at https://github.com/microsoft/JARVIS/tree/main/easytool. The inference setting of Yuan et al. (2024) differs from ours as we follow the official inference prompts provided with StableToolBench.. Additionally, we compare against DRAFT (Qu et al., 2025), a concurrent work, which optimizes tool documentation only in an iterative manner  [^4] We use the optimized documentation very recently open-sourced at https://github.com/quchangle1/DRAFT. Note that  Qu et al. (2025) tested only on the I2-Cat and I3-Inst subsets.. An evaluation LLM is used to judge whether a response adequately answers a user query, due the dataset’s free-form output design. We follow the official pipeline, using official ReAct prompts and report solvable pass rate, which measures the percentage of queries deemed solvable by the evaluation LLM. Further details are provided in appendix B.
Optimization Details.
Play2Prompt first uses beam search to optimize tool-use examples. We set $\lambda=1$, $N_{E}=3$, limit depth to $3$, and explore $L=3$ beams per node. We use beam width $W=10$ for BFCL and $W=3$ for StableToolBench to generate and select the top $W$ examples for each tool, which then are passed to the documentation optimization procedure. Beam search is again applied to select the best tool documentation. We employ llama-3.1-8b-instruct as both ${\mathcal{G}}_{E}$ and ${\mathcal{G}}_{D}$.
[…]
Figure 2: Average (weighted) accuracy improvements with Play2Prompt on BFCL, for different parameter description dropout $p$.
3.2 Results and Analyses
Results on BFCL and StableToolbench.
In table 1, we summarize the results on BFCL with LLaMA and GPT task models, both with and without tool-use examples and tool documentation produced by Play2Prompt. With Play2Prompt, absolute gains of 4-7% are observed for the open models and GPT-3.5, while GPT-4o achieves 3% increase in average accuracy despite its already high baseline. Notably, Play2Prompt addresses challenging categories such as REST for LLaMA-8B and Hammer-7B, and Multiple-Parallel for all models, yielding improvements of 10-17%. These gains suggest that optimized in-context examples and documentation can correct specific shortcomings in tool usage. Moreover, for the more difficult multi-tool queries in Multiple and Multiple-Parallel, we see larger gains compared to single-tool, even when we optimize each tool independently.
In table 2, we show the solvable pass rates on StableToolBench for LLaMA and GPT models with and without Play2Prompt. Our approach surpasses all baselines with average gains of 5-7% across LLaMA models and GPT-3.5, and outperforms the specifically designed EasyTool, which is not fully zero-shot as it requires in-context demonstrations during optimization. Play2Prompt is also consistent across subsets, avoiding large performance drops in multiple subsets when using EasyTool, highlighting the benefits of real-time “tool-play” and evaluation of candidate examples and documentation during optimization. Moreover, as GPT-4o exhibits more sensitivity to documentation and in-context demonstration changes on this dataset, EasyTool degrades GPT-4o on all subsets, whereas Play2Prompt achieves a 2% improvement on average. Overall, Play2Prompt essentially boosts performance of models up to the baseline performance of the larger models. Compared to DRAFT, which optimizes tool documentation only, Play2Prompt outperforms for all four models on both I2-Cat and I3-Inst. This suggests that optimizing not only documentation, as DRAFT does, but also generating tool-use examples, can synergize and achieve even larger improvements. Additionally, our optimization model of LLaMA-8B is much smaller than the GPT-4o model used in DRAFT, further confirming the effectiveness of Play2Prompt. It is noteworthy that most test samples for all 6 subsets in StableToolBench, as well as the Multiple and Multiple-Parallel subsets in BFCL, are multi-tool queries (see Section 3.1), underscoring the effectiveness of Play2Prompt, which operates on single tools independently during optimization.
@@ 3.2 Results and Analyses
To evaluate the robustness of Play2Prompt under incomplete documentation, we simulate noisy tool descriptions on the BFCL Executable dataset by randomly dropping parameter descriptions with increasing probabilities $p\in\{0.0,0.5,1.0\}$, while retaining overall tool descriptions and parameter names. This setup reflects varying degrees of real-world documentation sparsity. As expected, the degradation of documentation leads to reduced baseline accuracy across all models.
Despite these challenges, Play2Prompt consistently improves tool-use performance across all models and noise levels. As shown in figure 2, Play2Prompt achieves steady accuracy gains, with average improvements growing from 5.0% at $p=0.0$ and reaching 7.1% at $p=1.0$. Notably, the benefit of Play2Prompt becomes even more pronounced as documentation quality deteriorates, demonstrating its ability to compensate for missing parameter details and its strong robustness in low-resource settings. Detailed per-model results can be found in appendix C.
@@
Ablation on Example Difficulty.
Search strategy Pass Rate # Proposals Explored
ReAct 64.3 -
+P2P-MC (depth$=1$) 65.2 3
+P2P-MC (depth$=5$) 66.7 15
+P2P-MC (depth$=5$) & $N_{E}$$=3$ 68.9 45
+P2P-Beam (depth$=5$) & $N_{E}$$=3$ 70.7 135
@@
We investigate alternative strategies for exploring the example-documentation space by comparing beam search in Play2Prompt to Monte Carlo search (MC) at various depths, sharing the same sample proposal scheme but with different number of proposals and sub-sampling approach. We also study the role of $N_{E}$ self-reflection steps during example generation. Results are shown in table 5, focusing on I3-Inst, which combines multiple tools from different categories and is thus more challenging. The results show that MC underperforms beam search due to its limited exploration and is more subject to randomness, while deepening the search and adding self-reflection actions significantly improves performance, highlighting the importance of systematic search in optimizing examples and documentation. Given the tradeoff between performance and efficiency, a suitable search method can be chosen when using Play2Prompt based on the user’s computational budget.
@@
Dataset Completeness Conciseness Accuracy
STB P2P Raw Equal P2P Raw Equal P2P Raw Equal
85% 2% 13% 35% 47% 18% 43% 19% 37%
BFCL P2P Raw Equal P2P Raw Equal P2P Raw Equal
79% 5% 17% 15% 70% 15% 18% 19% 63%
Table 6: Human evaluation on tool documentation quality, comparing completeness, conciseness, and accuracy of Play2Prompt versus raw (initial) documentation.
Examples Naturalness Alignment
BFCL Test Set 4.39 4.68
Play2Prompt 4.17 4.52
Table 7: Human evaluation on tool usage example (question-function call) quality, rated on a scale of 1 (low) to 5 (high).
Human Evaluation of Generation Quality.
Since our primary objective is to improve LLMs’ tool-use capabilities, our main evaluation focuses on tool-use accuracy, which directly measures the effectiveness of our approach. While the quality of documentation and examples may not directly reflect tool-use improvements, they can provide complementary insights—particularly in enhancing clarity for human users and serving as secondary evidence of optimization quality. To assess these aspects, we conducted human evaluations on 50 randomly selected documentation outputs and 30 tool-use examples using six annotators. For documentation, we follow Qu et al. (2025) and assess completeness, conciseness, and accuracy relative to the original dataset-provided descriptions. For tool-use examples, we adopt the evaluation criteria from Shen et al. (2024), scoring the naturalness of the query and its alignment with the function call [^5] We omit the complexity metric used in Shen et al. (2024), as Play2Prompt generates single-tool usage examples., each on a scale of 1 to 5. As shown in table 7, our optimized documentation achieves significantly higher completeness and comparable or better accuracy, though with reduced conciseness due to increased verbosity. As shown in table 7, the zero-shot-generated examples are rated similarly to manually curated ones in terms of both naturalness and alignment, demonstrating their practical utility.
@@ 4 Related Work
LLMs for Tool Use.
Recent years have seen notable advances in employing large language models (LLMs) as agents to master tool use for solving complex tasks (Mialon et al., 2023; Qin et al., 2024a), thereby enhancing LLMs’ capabilities in multi-modal understanding (Gupta and Kembhavi, 2023; Surís et al., 2023; Wu et al., 2023), programming tools (Gao et al., 2023; Paranjape et al., 2023; Team et al., 2023; Zhang et al., 2023b; Cai et al., 2024), and other domain-specific functionalities. The conventional approach involves training base models with tool-use data (Thoppilan et al., 2022; Dubey et al., 2024) or fine-tuning LLMs (Patil et al., 2023; Schick et al., 2023; Yang et al., 2023; Parisi et al., 2022; Liu et al., 2025; Lin et al., 2025), but may require continual learning as new tools are added. Hao et al. (2023) addressed scalability by training tool embeddings plug-and-play usage, though still requiring labeled data. Alternatively, LLMs can be augmented with meta-prompts or tool-use instructions at inference time (Lu et al., 2023; Shen et al., 2023; Song et al., 2023; Qin et al., 2024b; Zhuang et al., 2024). As the range of applications and tools expands, enhancing LLMs’ capacity to handle new tools remain pivotal, which we address through Play2Prompt.
Tool-Use Instructions and Optimization.
Tool documentation and example demonstrations are key to prompting LLMs for effective tool use. Hsieh et al. (2023) highlighted the risk of hallucinations when documentation is lacking, and Xu et al. (2023) showed performance declines without in-context examples. To automate generating tool-use instances, Shen et al. (2024) leveraged a graph of tool relations for back-instructing queries, relying on the availability of external tool graphs. Yuan et al. (2024) proposed direct prompting to rewrite tool documentation, which relies on labeled documentation examples and lacks systematic optimization and the ability to measure optimization quality. Qu et al. (2025) explored LLMs interacting with tools, focusing on tool documentation only with single-thread iterative refinement. Automatic prompt tuning (Pryzant et al., 2023; Wang et al., 2024) adapts prompts to domain-specific tasks but require held-out test sets, rendering it unsuitable for zero-shot tool instruction rewriting (Wu et al., 2024). These constraints underscore the need for approaches that optimize tool instructions and demonstrations without labeled data or manual effort, which Play2Prompt achieves by interacting directly with the tool itself.
@@ 5 Conclusion
Acknowledgments
This research is supported by the MIT-IBM Watson AI Lab and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd. under the Innovation and Technology Commission’s InnoHK Scheme. The views and conclusions are those of the authors and should not be interpreted as representing those of IBM.
@@ Acknowledgments
In this work, Play2Prompt optimizes a single tool, but it’s still applicable to queries that require multiple tools by optimizing each tool independently, and then using the optimized examples and documentation from multiple tools together at inference. Although results (on Multiple and Multiple-Parallel in BFCL and the entirety of StableToolBench) on multi-tool queries show sizeable performance gains, scaling the optimization process itself from single-tool scenarios to multiple tools can likely enhance LLM’s tool use effectiveness even more. Additionally, for example demonstrations, we use rejection sampling to generate tool invocations first, which do not work for functions whose parameter space is too large, for instance parameters that take long ID string inputs or authentication tokens that require calls to other tools beforehand. Exploring multi-tool dependencies could potentially resolve this issue and improve tool play. In our work we focus on tool documentation and demonstrations only, relegating other information as meta-information, which could be potential next steps to explore.
@@ Appendix A Dataset License
Appendix B Additional Implementation Details
For task LLM ${\mathcal{M}}_{T}$, llama-3.1-8b-instruct is used for LLaMA-8B, llama-3.3-70b-instruct for LLaMA-70B, gpt-3.5-turbo-0125 for GPT-3.5, gpt-4o-2024-11-20 for GPT-4o, toolace-2-llama-3.1-8b for ToolACE-8B, and hammer2.1-7b for Hammer-7B. On StableToolBench, we use llama-3.3-70b-instruct as the evaluation LLM, which produces stable assessments.
For GPT models, we keep the exact same inference setting as used in the original benchmark, except for requiring it to return a single action at each step from the provided toolset by supplying a flag to the OpenAI API. For LLaMA models, on BFCL we follow the exact official inference prompts, and on StableToolBench we adapt the ReAct prompts into LLaMA-3’s prompt format but keep everything else fixed as much as possible. For ToolACE and Hammer on BFCL, we also follow the exact official inference prompts.